Graph 6y+5=-2x+5. Also, give the x-intercept, the y-intercept and the slope.
Answer by mananth(16946)

2 x + 6 y = 0
y= - 1/ 3 x + 0

Assume values of x and plug in the equation to get values of y.
Plot the points
x 0 3 -3 6
y= 0 -1 1 -2
(x,y) ( 0 , 0 ),( 3 , -1 ),( -3 1 ),(6,-2)

Slope = -1/3